Hi can someone please help me.
How high is a tree that casts a 35ft shadow at the same time a 44​ft post casts a shadow which is 12ft ​long?

Answer by rothauserc(4718)

Use similar right triangles, then
:
let AB = height of tree and AC = 35 is the length of the tree's shadow
:
tangent of angle BCA = 44/12 = 3.67 and
:
inverse tangent of BCA = 74.74 degrees, then
:
tan 74.74 = AB / 35, therefore
:
AB = 3.67 * 35 = 128.45
:
****************************
The tree is 128.45 feet high
****************************
: